OpenStudy (anonymous):

which of the following could be a ratio between the lengths of the two legs of a 30-60-90 triangle?

OpenStudy (anonymous):

the following?

OpenStudy (anonymous):

one should have \(\sqrt 2\) or something

OpenStudy (anonymous):

it's \(1,\frac 12 ,\frac{\sqrt 2}2\)

OpenStudy (wolf1728):

Here's a web page about 30° 60° 90° triangles: http://1728.org/trig2.htm The possible ratios could be: 2; sqrt(3); 1; (2/sqrt(3)); (1/sqrt(3)); sqrt(3) /2 ; or a half